Given that C=C1 = cy and C1=500, c3/4, y=1000
1)total consumption is 500
2)total consumption is 1250
3)aggregate demand is 1000
4)marginal propensity to consume is 1/4
1
Expert's answer
2019-09-18T09:41:39-0400

Given that C = C1 + cy, C1 = 500, c = 3/4, and y = 1000 the correct answer are 2 and 3, because C = 500 + 3/4*1000 = 1250 and AD = y = 1000.